OpenStudy (anonymous):

Triangle A B C is shown with horizontal base B C 8 feet, one vertical side C A 6 feet, and the remaining side A B an unknown number of feet, indicated with a question mark. Triangle J K L is similar to triangle A B C. Side K L corresponds to B C and is twelve feet long. Side J L corresponds to A C and is 9 feet long. Side J K corresponds to A B and is fifteen feet long. What is the length of side A B?

OpenStudy (mayankdevnani):

Where is triangle ABC and JKL ?? @twilaswift

OpenStudy (mayankdevnani):

always remember that in similar triangles we have, \[\huge \bf \frac{AC}{JL}=\frac{BC}{KL}=\frac{AB}{JK}\]

OpenStudy (anonymous):

Okay... this is the answers I have: 10 feet 11 feet 12 feet 13 feet I think it's 12 feet.

OpenStudy (mayankdevnani):

\[\huge \bf \frac{6}{9}=\frac{8}{12}=\frac{AB}{15}\]

OpenStudy (mayankdevnani):

solve for AB? @twilaswift

OpenStudy (anonymous):

13.

OpenStudy (anonymous):

No wait it's 11.

OpenStudy (mayankdevnani):

\[\huge \bf \frac{6}{9}=\frac{2}{3}\] so, \[\huge \bf \frac{2}{3}=\frac{AB}{15}\] so solve for x?

OpenStudy (mayankdevnani):

@twilaswift

OpenStudy (anonymous):

12

OpenStudy (mayankdevnani):

do cross multiply, we get \[\huge \bf 15 \times 2=AB \times 3\] divide 3 from both the sides, we get \[\huge \bf AB=\frac{15 \times 2}{3}\]

OpenStudy (mayankdevnani):

can you solve it? @twilaswift

OpenStudy (anonymous):

30?

OpenStudy (anonymous):

But my answers are only 10, 11, 12 and 13 no 30.

OpenStudy (mayankdevnani):

\[\huge \bf 15 \times 2=30\] then divide 30 by 3

OpenStudy (mayankdevnani):

so what is our answer? @twi

OpenStudy (mayankdevnani):

@twilaswift